OpenStudy (anonymous):

6x - 2y = 5 3x - y = 10 Solve the system of equations.

OpenStudy (anonymous):

I'd solve the second equation for y and substitute in the first equation

OpenStudy (anonymous):

\[6x-2(3x-10)=5\] and solve for x

OpenStudy (anonymous):

Dividing the top eq by 2 would give 3x-y=5/2

hero (hero):

Divide the first equation by 2 and realize that the lines are parallel

OpenStudy (anonymous):

There is no solution possible--they are parallel lines

OpenStudy (anonymous):

thank you!